# Shaping Dynamics Through Memory: A Study of Reservoir Profiles in Open Quantum Systems

### v1

@@ I Introduction
Open quantum systems provide a natural framework for describing the dynamics of physical platforms that interact with external environments. In photonic and condensed-matter settings, such environments are often modeled as reservoirs coupled to a localized mode or guided field [Ferreira2021, Longhi2006]. These interactions allow for the exchange of energy and information, leading to complex non-Markovian dynamics and quantum correlations [Budini2022, Berrada2025]. The properties of the reservoir as its spectral density, spatial correlations, and characteristic decay, play a decisive role in shaping system evolution [Galve2016, Gaidi2026]. Depending on the structure of these correlations, the dynamics may range from fully Markovian, where memory effects are negligible, to strongly non-Markovian regimes in which information can effectively flow back from the environment to the system. Understanding how different reservoir profiles influence observable quantities is therefore essential for both fundamental studies and technological applications such as light transport, dissipative engineering, and quantum information processing [Bylicka2014, Li2022, Shirai2021, ElAllati2024].
In waveguide–reservoir architectures, environmental memory manifests itself explicitly through convolution equations that couple the field amplitude at a given position to its entire spatial interaction history [Vega2017, Tufarelli2014]. This nonlocal dependence highlights how reservoir structure governs spatial propagation rather than simply modifying attenuation strengths. Different reservoir kernels capture distinct physical mechanisms: Markovian reservoirs correspond to memoryless exponential correlations and serve as a reference model for irreversible propagation [Gardiner1985]; Lorentzian reservoirs encode finite spatial correlation lengths associated with cavity-filtered or resonant spectral densities [Garraway1997]; Gaussian and Uniform kernels describe structured or broadband environments, often relevant in engineered photonic media or environments shaped by fabrication constraints [Vasile2011, Calajo2019]. Although each of these kernels has been studied individually in various contexts, a systematic comparison of their impact on propagation within a single, controlled scenario remains largely unexplored [Breuer2016, Liu2011].
@@ II Hamiltonian and Heisenberg equations solution
Here, $\alpha(\beta_{n})$ represents the spectral coupling function, encoding how strongly the guided mode interacts with each mode of the environment. The interaction term $r_{n}^{\dagger}a+r_{n}a^{\dagger}$ facilitates the coherent exchange of energy while discarding rapidly oscillating counter-rotating terms [bloch1940magnetic, jaynes2005comparison]. Consequently, the nature of the reservoir, specifically the spectral density defined by the distribution of $\alpha(\beta_{n})$, determines the dynamical regime, which can induce significant non-Markovian features.
@@
as indicated by the Wiener-Khichin theorem [goodman2015statistical]. To evaluate the different possible regimes generated by each distribution, consider Table 1 with the particular choice of three probability distributions with their autocorrelations and Full Widths at Half Maximum (FWHM), the latter being the key parameter to match the communication window between the photons propagating in the waveguides and the reservoir modes.
@@
$q.n.$ in equation (6) refers to quantum noise terms [clerk2010introduction], which can be neglected when choosing states where the reservoir starts the dynamics empty.
@@ III Transmission
Figure 2: Transmission with $\gamma=2\alpha$ for three reservoirs highlighted with a solid line and the Hermitian situation where $\gamma=0$ with a dashed line.
@@
The results obtained for $|f|^{2}$ are consistent with Fermi’s Golden Rule [fox2006quantum], which serves as the macroscopic limit of the problem. In this Markovian regime, the decay rate is directly proportional to the spectral density at resonance, given by $\Gamma\propto 2\pi\alpha^{2}\rho(\beta)$. By comparing the distributions for a fixed Full Width at Half Maximum (FWHM), we observe that the Uniform distribution yields the most pronounced decay. This occurs because the rectangular profile concentrates the spectral weight entirely within the bandwidth, maximizing $\rho(\beta_{c})$. It is followed by the Gaussian case, while the Lorentzian distribution exhibits the slowest decay. The latter behaves this way because its heavy tails distribute a significant portion of the spectral density away from the resonance, resulting in the lowest peak value $\rho(\beta_{c})$ among the three configurations.
@@ IV Transparence
Optical transparency [guo2009observation] refers to the photon transport capability within a waveguide without significant absorption by the reservoir. In this context, Loss-Induced Transparency (LIT) [beder2024quantum, silva2025anomalous] stands out as a counter-intuitive non-Hermitian effect, where enhancing dissipation in specific channels favors overall transmission. While LIT is traditionally associated with spatial losses, we present in Figure 4 a dynamical analogue of this effect: transparency assisted by memory suppression.
@@ V Non-Markovianity measure
To quantify the memory effects of the dynamics, we employ the non-Markovianity measure proposed by Breuer, Laine, and Piilo (BLP measure) [chiriaco2025computable, guarnieri2014quantum, breuer2009measure], which is based on the variation of distinguishability between quantum states. Although the general definition of this measure requires an optimization over all possible pairs of initial states to detect information backflow, the system studied here, a waveguide in the single-excitation regime interacting with the environment, belongs to the class of amplitude damping channels. For this specific class, it has been analytically demonstrated that the pair of states maximizing the information flow is the orthogonal pair formed by the vacuum state $|0\rangle$ and the excited state $|1\rangle$ [breuer2009measure, breuer2016colloquium]. Consequently, the trace distance becomes exactly equal to the absolute value of the photon amplitude function, $|f(z)|$. Thus, the non-Markovianity measure $\mathcal{N}$ simplifies to the integral of the positive derivative of $|f(z)|$,
@@
In the Hermitian limit ($\gamma=0$), where the reservoir effectively behaves as a secondary waveguide, the BLP measure yields $\mathcal{N}_{0}=\lfloor\alpha L/\pi\rfloor+[1-\Theta(\tan(\alpha L))]|\cos(\alpha L)|$ for a waveguide of length $L$. In the simulation shown in Figure 5, we used $L=100\pi$. All curves in Fig. 5 exhibit a monotonic decay. Regarding memory effects, we observe that spectral distributions which are more concentrated and possess lighter tails correspond to stronger non-Markovian features in the system-reservoir interaction.
The physical interpretation of the results obtained via the BLP measure can be further elucidated through the pseudo-mode approach [Garraway1997, Mazzola2009, Pleasance2020], adapted here to the spatial domain. In this framework, the reservoir’s density of propagation constants is analytically mapped onto a set of fictitious auxiliary modes interacting with the main waveguide. A Lorentzian distribution of propagation constants, characterized by long, smooth tails in reciprocal space, mathematically corresponds to a single pole in the complex plane. This is equivalent to coupling the waveguide to a single lossy pseudo-mode, a simple structure that favors purely exponential spatial decay, resulting in essentially Markovian dynamics along the $z$-axis.Conversely, distributions with compact support, such as the uniform distribution, cannot be represented by a single simple pole. The presence of sharp cutoffs in the allowed propagation constants introduces singularities that effectively require a superposition of multiple pseudo-modes to describe the spatial interaction. The BLP measure thus captures the interference effects between these modes during propagation, creating a coherent framework that allows to evaluate the revivals of photon amplitude on the main waveguide at subsequent positions, characterizing spatial information backflow.

### v2

# Non-markovian reservoir profile effects on dynamics of light within a single waveguide

@@ I Introduction
Open quantum systems provide a natural framework for describing the dynamics of physical platforms that interact with external environments. In photonic and condensed-matter settings, such environments are often modeled as reservoirs coupled to a localized mode or guided field [14, 26]. These interactions allow for the exchange of energy and information, leading to complex non-Markovian dynamics and quantum correlations [7, 2]. The properties of the reservoir as its spectral density, spatial correlations, and characteristic decay, play a decisive role in shaping system evolution [17, 16]. Depending on the structure of these correlations, the dynamics may range from fully Markovian, where memory effects are negligible, to strongly non-Markovian regimes in which information can effectively flow back from the environment to the system. Understanding how different reservoir profiles influence observable quantities is therefore essential for both fundamental studies and technological applications such as light transport, dissipative engineering, and quantum information processing [8, 24, 29, 13].
In waveguide–reservoir architectures, environmental memory manifests itself explicitly through convolution equations that couple the field amplitude at a given position to its entire spatial interaction history [33, 31]. This nonlocal dependence highlights how reservoir structure governs spatial propagation rather than simply modifying attenuation strengths. Different reservoir kernels capture distinct physical mechanisms: Markovian reservoirs correspond to memoryless exponential correlations and serve as a reference model for irreversible propagation [18]; Lorentzian reservoirs encode finite spatial correlation lengths associated with cavity-filtered or resonant spectral densities [19]; Gaussian and Uniform kernels describe structured or broadband environments, often relevant in engineered photonic media or environments shaped by fabrication constraints [32, 9]. Although each of these kernels has been studied individually in various contexts, a systematic comparison of their impact on propagation within a single, controlled scenario remains largely unexplored [4, 25].
@@ II Hamiltonian and Heisenberg equations solution
Here, $\alpha(\beta_{n})$ represents the spectral coupling function, encoding how strongly the guided mode interacts with each mode of the environment. The interaction term $r_{n}^{\dagger}a+r_{n}a^{\dagger}$ facilitates the coherent exchange of energy while discarding rapidly oscillating counter-rotating terms [3, 23]. Consequently, the nature of the reservoir, specifically the spectral density defined by the distribution of $\alpha(\beta_{n})$, determines the dynamical regime, which can induce significant non-Markovian features.
@@
as indicated by the Wiener-Khichin theorem [20]. To evaluate the different possible regimes generated by each distribution, consider Table 1 with the particular choice of three probability distributions with their autocorrelations and Full Widths at Half Maximum (FWHM), the latter being the key parameter to match the communication window between the photons propagating in the waveguides and the reservoir modes.
@@
$q.n.$ in equation (6) refers to quantum noise terms [11], which can be neglected when choosing states where the reservoir starts the dynamics empty.
@@ III Transmission
Figure 2: Transmission with $\gamma=2\alpha$ for three reservoirs highlighted with solid and dashed lines and the Hermitian situation where $\gamma=0$ with a dotted line.
@@
The results obtained for $|f|^{2}$ are consistent with Fermi’s Golden Rule [15], which serves as the macroscopic limit of the problem. In this Markovian regime, the decay rate is directly proportional to the spectral density at resonance, given by $\Gamma\propto 2\pi\alpha^{2}\rho(\beta)$. By comparing the distributions for a fixed Full Width at Half Maximum (FWHM), we observe that the Uniform distribution yields the most pronounced decay. This occurs because the rectangular profile concentrates the spectral weight entirely within the bandwidth, maximizing $\rho(\beta_{c})$. It is followed by the Gaussian case, while the Lorentzian distribution exhibits the slowest decay. The latter behaves this way because its heavy tails distribute a significant portion of the spectral density away from the resonance, resulting in the lowest peak value $\rho(\beta_{c})$ among the three configurations.
@@ IV Transparence
Optical transparency [22] refers to the photon transport capability within a waveguide without significant absorption by the reservoir. In this context, Loss-Induced Transparency (LIT) [1, 30] stands out as a counter-intuitive non-Hermitian effect, where enhancing dissipation in specific channels favors overall transmission. While LIT is traditionally associated with spatial losses, we present in Figure 4 a dynamical analogue of this effect: transparency assisted by memory suppression.
@@ V Non-Markovianity measure
To quantify the memory effects of the dynamics, we employ the non-Markovianity measure proposed by Breuer, Laine, and Piilo (BLP measure) [10, 21, 6], which is based on the variation of distinguishability between quantum states. Although the general definition of this measure requires an optimization over all possible pairs of initial states to detect information backflow, the system studied here, a waveguide in the single-excitation regime interacting with the environment, belongs to the class of amplitude damping channels. For this specific class, it has been analytically demonstrated that the pair of states maximizing the information flow is the orthogonal pair formed by the vacuum state $|0\rangle$ and the excited state $|1\rangle$ [6, 5]. Consequently, the trace distance becomes exactly equal to the absolute value of the photon amplitude function, $|f(z)|$. Thus, the non-Markovianity measure $\mathcal{N}$ simplifies to the integral of the positive derivative of $|f(z)|$,
@@
In the Hermitian limit ($\gamma=0$), where the reservoir effectively behaves as a secondary waveguide, the BLP measure yields $\mathcal{N}_{0}=\lfloor\alpha L/\pi\rfloor+[1-\Theta(\tan(\alpha L))]|\cos(\alpha L)|$ for a waveguide of length $L$. In the simulation shown in Figure 5, we used $\alpha L=100\pi$. All curves in Fig. 5 exhibit a monotonic decay. Regarding memory effects, we observe that spectral distributions which are more concentrated and possess lighter tails correspond to stronger non-Markovian features in the system-reservoir interaction.
The physical interpretation of the results obtained via the BLP measure can be further elucidated through the pseudo-mode approach [19, 27, 28], adapted here to the spatial domain. In this framework, the reservoir’s density of propagation constants is analytically mapped onto a set of fictitious auxiliary modes interacting with the main waveguide. A Lorentzian distribution of propagation constants, characterized by long, smooth tails in reciprocal space, mathematically corresponds to a single pole in the complex plane. This is equivalent to coupling the waveguide to a single lossy pseudo-mode, a simple structure that favors purely exponential spatial decay, resulting in essentially Markovian dynamics along the $z$-axis.Conversely, distributions with compact support, such as the uniform distribution, cannot be represented by a single simple pole. The presence of sharp cutoffs in the allowed propagation constants introduces singularities that effectively require a superposition of multiple pseudo-modes to describe the spatial interaction. The BLP measure thus captures the interference effects between these modes during propagation, creating a coherent framework that allows to evaluate the revivals of photon amplitude on the main waveguide at subsequent positions, characterizing spatial information backflow.